OpenStudy (anonymous):

Take any number (except for 1). Square that number and then subtract one. Divide by one less than your original number. Now subtract your original number. Did you reached 1 for an answer? You should have. How does this number game work? Hint: Redo the number game using a variable instead of an actual number and rewrite the problem as one rational expression. How did the number game use the skill of simplifying rational expressions? Create your own number game using the rules of algebra and post it for your classmates to solve. Think about values that may not work. State whether you

OpenStudy (anonymous):

(x^2 - 1) / (x-1) - x = x+1 - x = 1 own game: think of a number from 1 to 9 add 1 to it multiply your answer by 3 add 1 to the result now add first number you thought of to your answer ask what the answer is the answer will be a multiple of 4 to find the original number subtract 4 from the answer and divide by 3. this works because of the following identity: (x+1)*3 +1 + x = 4x + 4

OpenStudy (anonymous):

are u happy with that korie?

OpenStudy (anonymous):

yes but can you give me one more example please